Shift summary for Day, 8/2/99

roland winston (r_winston@yahoo.com)
Mon, 2 Aug 1999 13:23:44 -0700 (PDT)

Date: Mon, 02 Aug 1999 06:50:40 -0500
From:Roland Winston <roland@hep.UCHICAGO.edu>
To: shiftsum@vivsgi.fnal.gov
Subject: Shift summary for Day, 8/2/99


Shift Summary for Day Shift, 08/2/99

On Shift: NS,RW, Ryan Morfin (UofC undergrad visiting shift person)

Hours of beam available at all intensities: 4.6
Hours of beam available at requested intensity: 4.6
Hours of E832 data (special runs not included): 4.6
Hours of special runs (laser scans, muon runs...): 1.2

Accounting of lost time: Tevatron off for about 3.5 hrs., plus .4 for
various lost spills.
========================================================================
Runs ended this shift:
Run 14085 beam 832 run
14088 laser scan
14089 muon run

still running at end of shift 14090

+++++++++++++++++++++++++++++++++++++++++++++++++++++++++
Tapes taken to Feynmann center: QKa20x (a=C,S,K and x=7,8,9) and QKB071

=======================================================================
Accesses during shift: None
========================================================================
Overview of shift: Waiting for Tevatron to return to operation, plus
setting up 2e12 run with unlimited ke3s.

========================================================================
Description of Shift:

run spill time comment

------------------------------------------------------------------------
14085 07:00 Day shift takes over, data taking in progress
106 ktevana errors, ran Kkfrend
109 Automatic Fped ran
123 Ktsad.exe not runningg on atev2
160 Beam Off MI Q847 problems, experts
investigating.
180 09:00 Beam returns
187 Beam Off, MCR says is it main injerctor RF vacuum problems.
14085 212 09:42 End run (beam still off)
14088 Laser scan
11:07 Run down the magnet for doing a Muon run
11:55 Laser run ended, called MCR to set up beamline
for a muon run.
14089 12:00 Muon run started, waiting for beam.
1 12:01 Muons arrive from accelerator.
3 muon run ended, 200000 events.
14090 12:10 NM4AN up, MCR put sweepers back to data mode
and started init for special low intensity
2e12 data mode with unlimited ke3s.
12:13 onsplit asked for qkmu11 be mounted in ktev4
14090 6 12:25 first Spill with beam.
7 Val warning, Laser RMS too wide channel 1895.
87 Val warning, Laser RMS too wide channel 1557.
90 Beam gone.
96 beam is back!
121 15:00 Ktev ana errors, we ran Kfrend.



========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift
Keep going for about 3 days with the special
2e12 run with unlimited ke3s.










===
Roland Winston
Professor, Dept. of Physics
The University of Chicago
5640 South Ellis Avenue
Chicago, IL 60637
(773) 702-7756

_____________________________________________________________
Do You Yahoo!?
Free instant messaging and more at http://messenger.yahoo.com